OpenStudy (anonymous):

y^2-y-2=0 Complete the square.

OpenStudy (anonymous):

[y-(1/2)]^2-9/4

OpenStudy (anonymous):

whats the soluution set

OpenStudy (anonymous):

(y-2)(y+1)=0 y-2=0 y=2 y+1=0 y=-1